# Stochastic Optimal Control of an Industrial Power-to-Heat System with High-Temperature Heat Pump and Thermal Energy Storage

### v2

@@ H.1 Construction of Running Costs
with $I_{1}=\int_{\mathbb{R}}\tfrac{\Sigma_{S}v+m^{S}}{\sqrt{2\pi}\sqrt{1-\rho^{2}}}\exp\big{\{}-\tfrac{(v-\rho u)^{2}}{2(1-\rho^{2})}\big{\}}\mathrm{d}v$. For solving integral $I_{1}$ one has to observe that the function inside is the density of a random variable $V\sim\pazocal{N}(\rho u,1-\rho^{2})$ therefore
$I_{1}$ $=\int_{\mathbb{R}}\tfrac{\Sigma_{S}v}{\sqrt{2\pi}\sqrt{1-\rho^{2}}}\exp\big{\{}-\tfrac{(v-\rho u)^{2}}{2(1-\rho^{2})}\big{\}}\mathrm{d}v+\int_{\mathbb{R}}\tfrac{m^{S}}{\sqrt{2\pi}\sqrt{1-\rho^{2}}}\exp\big{\{}-\tfrac{(v-\rho u)^{2}}{2(1-\rho^{2})}\big{\}}\mathrm{d}v$ (232)
@@
Going back to the calculation of $E(t,k,R)$ and substituting the above we get
$E(t,k,R)$ $=\int_{u(a)}^{u(b)}\tfrac{1}{\sqrt{2\pi}}e^{k(\Sigma_{W}u+m^{W})}e^{-\tfrac{u^{2}}{2}}(\Sigma_{S}\rho u+m^{S})\mathrm{d}u$ (234)
$=\int_{u(a)}^{u(b)}\tfrac{\Sigma_{S}\rho u+m^{S}}{\sqrt{2\pi}}\exp\big{\{}-\tfrac{u^{2}-2k\Sigma_{W}u+(k\Sigma_{W})^{2}-(k\Sigma_{W})^{2}}{2}+km^{W}\big{\}}\mathrm{d}u$ (235)
$=\int_{u(a)}^{u(b)}\tfrac{\Sigma_{S}\rho u+m^{S}}{\sqrt{2\pi}}\exp\big{\{}-\tfrac{(u-k\Sigma_{W})^{2}}{2}+\tfrac{(k\Sigma_{W})^{2}}{2}+km^{W}\big{\}}\mathrm{d}u$ (236)
@@
with $I_{2}=\int_{u(a)}^{u(b)}\tfrac{\Sigma_{S}\rho u+m^{S}}{\sqrt{2\pi}}\exp\big{\{}-\tfrac{(u-k\Sigma_{W})^{2}}{2}\big{\}}\mathrm{d}u$. Analogous to $I_{1}$, the function inside $I_{2}$ is the density of the random variable $U\sim\pazocal{N}(k\Sigma_{W},1)$ which yields
$I_{2}$ $=\Sigma_{S}\rho\int_{u(a)}^{u(b)}\tfrac{u}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u+m^{S}\int_{u(a)}^{u(b)}\tfrac{1}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u$ (238)
@@
$I_{3}$ $=\Sigma_{S}\rho\int_{u(a)}^{u(b)}\tfrac{u-k\Sigma_{W}+k\Sigma_{W}}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u$ (240)
$=\Sigma_{S}\rho\bigg{(}\int_{u(a)}^{u(b)}\tfrac{u-k\Sigma_{W}}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u+k\Sigma_{W}\int_{u(a)}^{u(b)}\tfrac{1}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u\bigg{)}$ (241)
$=\Sigma_{S}\rho\bigg{(}\tfrac{1}{\sqrt{2\pi}}(e^{-\tfrac{(u(a)-k\Sigma_{W})^{2}}{2}}-e^{-\tfrac{(u(b)-k\Sigma_{W})^{2}}{2}})+k\Sigma_{W}[\Phi(u(b)-k\Sigma_{W})-\Phi(u(a)-k\Sigma_{W})]\bigg{)}$ (242)
where $\int_{u(a)}^{u(b)}\tfrac{1}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u=\Phi(u(b)-k\Sigma_{W})-\Phi(u(a)-k\Sigma_{W})$. The last equality follows from the substitution $x=-\tfrac{(u-k\Sigma_{W})^{2}}{2}$ as well as solving the arising integral. Substitute the expression for $I_{3}$ into $I_{2}$ and hence back into $E(t,k,R)$ leads to
$E(t,k,R)=e^{\tfrac{(k\Sigma_{W})^{2}}{2}+km^{W}(t,y_{W})}\bigg{[}$ $\tfrac{\Sigma_{S}\rho}{\sqrt{2\pi}}(\exp\big{\{}-\tfrac{(u(a)-k\Sigma_{W})^{2}}{2}\big{\}}-\exp\{-\tfrac{(u(b)-k\Sigma_{W})^{2}}{2}\big{\}})$ (243)
$+(\Sigma_{S}\rho k\Sigma_{W}+m^{S}(t,y_{W},y_{S}))[\Phi(u(b)-k\Sigma_{W})-\Phi(u(a)-k\Sigma_{W})]\bigg{]}$ (244)
with $u(w)=\frac{\log(w)-m^{W}(t,y_{W})}{\Sigma_{W}}$. The closed-form expression for $E_{0}(t,k,R)$ is a derived from the expression of $E(t,k,R)$ by letting $m^{S}\equiv 1$ and $\Sigma_{S}\equiv 0$, i.e, price is constant with $S(t)=1$ for all $t\in[t_{n},t_{n+1})$.
@@
[…]
@@
with $\Delta t=t_{n+1}-t_{n}$ and weights $\omega_{l},~{}l=1,\ldots,k^{I}$. We find that $k^{I}=2$ is sufficient to get good results. Figure 13 shows the solution to $C_{n}^{2}$ obtained by an integral solver and the relative error induced by this Gauss-Legendre quadrature rule. It can be seen that the relative error is less than $2\%$, while acceleration the calculation by a factor of $100$ for the evaluation of the running cost.
Figure 13: Left: The cost functional $C_{n}^{2}$ calculated by an integral solver for $P^{\text{H}}_{n}=3500$ with respect to the wind speed and electricity price is shown. Right: The relative error between the Gauss-Legendre quadrature for $k^{I}=2$ and integral solver is shown.
H.2 Construction of Terminal Costs
In this work we consider terminal costs as in (32). This means, if the terminal TES temperature $R_{N}=r$ is below or above a critical temperature $r_{\text{crit}}\in[r_{\text{min}},r_{\text{max}}]$ a penalty or reward with respect to the difference $r_{\text{crit}}-r$ is applied. The penalty can be interpreted such that the TES has to be charged to the critical value and therefore additional costs are incurring. In converse, the reward is resulting from the liquidation of the saved energy. This is the energy that would be needed to charge the TES from $r_{crit}$ to the larger temperature $r$. In both cases, we assign a value to the energy associated with this difference given by $g_{\text{Ter}}$. These in turn are multiplied by predetermined fixed prices $s_{\text{Pen}},s_{\text{Liq}}\geq 0$, respectively, to evaluate the TES filling level. Usually the fixed prices satisfy $s_{\text{Pen}}\geq s_{\text{Liq}}$. In the following we will describe the penalization and liquidation process in order to construct $g_{\text{Ter}}$.
Penalization and Liquidation Cost
For the case $r_{\text{crit}}>r$ a penalization cost is applied, which is described by the following scenario. Since the TES temperature is below the critical threshold it needs to be charged manually. We assume that for all $t>t_{\text{E}}$ the produced wind power $P^{\text{W}}(t)=0$ and that the HTHP has to run on its maximum compressor shaft speed $d_{\text{max}}$. Therefore, the TES is charged as fast as possible and the energy consumption is covered solely by grid energy. From (13) the maximum energy consumption $P_{\text{max}}$ in this case is given by
@@ H.2 Construction of Terminal Costs
Moreover, the maximum compressor shaft speed $d_{\text{max}}$ yields the maximum outlet temperature
@@
The HTHP has to run on this continuous-time setting, until time $t_{\text{Ter}}=t_{\text{E}}+h(r,r_{crit})$ for which the TES temperature is raised to $r_{crit}$. Here, $h(r,r_{crit})>0$ describes the length of the charging period, which depends on the current TES state $r$ and the critical value $r_{crit}$. From the storage dynamics (3) we have
$r_{\text{crit}}=R(t_{\text{Pen}})=r+h(r,r_{crit})\frac{n_{\text{H}}\dot{m}c_{\text{F}}}{m_{\text{S}}c_{\text{S}}}(\tau^{\text{HT,out}}_{\text{max}}-T^{\text{SG,in}}),$ (248)
and determine the length of the charging period as
$h(r,r_{crit})=\frac{m_{\text{S}}c_{\text{S}}}{n_{\text{H}}\dot{m}c_{\text{F}}(\tau^{\text{HT,out}}_{\text{max}}-T^{\text{SG,in}})}(r_{\text{crit}}-r).$ (249)
The consumed electrical energy associated with $h$ is given by the function
$g_{\text{Ter}}(r,r_{\text{crit}})=h(r,r_{crit})P_{\text{max}}.$ (250)
The liquidation in the case $r_{\text{crit}}<r$ results from the saved energy, which it would require to charge the TES from the lower critical temperature $r_{\text{crit}}$ to the larger TES temperature $r$. Analogous to the penalization, we need to charge for a period of length $h(r_{crit},r)$ with the only difference that the role of $r$ and $r_{crit}$ have changed. Thus, the energy savings associated with the liquidation is given by
$g_{\text{Ter}}(r_{\text{crit}},r)=h(r_{crit},r)P_{\text{max}}.$ (251)
@@ Appendix I Details State Space Discretization
The state space $\pazocal{X}$ is going to be discretized into grid points, in order to calculate and store the value function for BDP. Moreover, we want to use the additive structure and closed-form solutions of the Ornstein-Uhlenbeck processes (16) to construct a family of discretizations $\pazocal{X}_{n},n=0,\ldots,N$. By this construction we make sure, that the value function is calculated in regions of interested, i.e. states that are located around the seasonalities. In the first step of the construction we define reference sets $\pazocal{X}^{Ref}_{n}\subset\pazocal{X}$
$\pazocal{X}^{Ref}_{n}=[r_{\text{min}},r_{\text{max}}]\times[\underline{w}_{n},\overline{w}_{n}]\times[\underline{s}_{n},\overline{s}_{n}],$ (253)
with boundary values $\underline{w}_{n},\overline{w}_{n}\in(0,\infty)$ and $\underline{s}_{n},\overline{s}_{n}\in(-\infty,\infty)$. For our specific case we construct these boundaries with the $k^{Ref}$-$\sigma$-rule motivated by the normal distribution of the closed-form solutions (27), leading to
$\underline{w}_{n}$ $=\exp(\mu_{W}(t_{n})-k^{Ref}\Sigma_{W}),~{}\overline{w}_{n}=\exp(\mu_{W}(t_{n})+k^{Ref}\Sigma_{W}),$ (254)
$\underline{s}_{n}$ $=\mu_{S}(t_{n})-k^{Ref}\Sigma_{S},~{}\overline{s}_{n}=\mu_{S}(t_{n})+k^{Ref}\Sigma_{S}.$ (255)
For states in $\pazocal{X}^{Ref}_{n}$ we want to have a good approximation, meaning that the errors introduced by the simplifications, namely discretization of states and actions as well as quantization of the expected value, should be small. The error of the state and action space discretization can be reduced by choosing a finer grid, i.e. increase the number of grid points. However handling the error introduced by the approximation of the expected value is more difficult. A main cause comes from the quantizer $\pazocal{Z}$. For $x\in\pazocal{X}^{Ref}_{n}$ the quantization points $\pazocal{T}_{n}(x,a,z)$, with $z\in\pazocal{Z}$, are unlikely to be contained in the discrete set $\widetilde{\pazocal{X}}_{n+1}$. Hence we need to interpolate or extrapolate the value function for these states and thus need to consider, not only interpolation but also extrapolation errors. In order to reduce extrapolation of the value function for states in the reference set, we enlarge them by choosing boundary values $w_{\min,n}\leq\underline{w}_{n},~{}\overline{w}_{n}\leq w_{\max,n}$ and $s_{\min,n}\leq\underline{s}_{n},~{}\overline{s}_{n}\leq s_{\max,n}$. Meaning that we construct an enlarged set $\pazocal{X}_{n}$ with $\pazocal{X}^{Ref}_{n}\subset\pazocal{X}_{n}$ given by
@@
in the sense that for (almost) every $x\in\pazocal{X}^{Ref}_{n-1}$, the transition $\pazocal{T}_{n-1}(x,a,z)$ will be contained in $\pazocal{X}_{n}$. Therefore, this helps to avoid or reduce the need of extrapolation for states in the reference set. The construction of the values for $w_{\min,n},w_{\max,n},s_{\min,n}$ and $s_{\max,n}$ is again motivated by the normal distribution of the the closed-form solution (27). For $w_{\min,n},w_{\max,n}$ we apply the $k^{Ext}$-$\sigma$-rule leading to
$w_{\min,n}=\exp(m^{W}_{n}(\underline{y}_{W,n-1})-k^{Ext}\Sigma_{W}),\quad w_{\max,n}=\exp(m^{W}_{n}(\overline{y}_{W,n-1})+k^{Ext}\Sigma_{W}),$ (257)
for $\underline{y}_{W,n-1}=\log(\underline{w}_{n-1})-\mu_{W}(t_{n-1})$ and $\overline{y}_{W,n-1}=\log(\overline{w}_{n-1})-\mu_{W}(t_{n-1})$. In order to construct $s_{\min,n},s_{\max,n}$ we proceed analogously. Considering that the electricity price and the wind speed are negatively correlated, i.e. prices are high if wind speeds are low and visa versa, we end up with the following values
$s_{\min,n}=m^{S}_{n}(\overline{y}_{W,n-1},\underline{y}_{S,n-1})-k^{Ext}\Sigma_{S},\quad s_{\max,n}=m^{S}_{n}(\underline{y}_{W,n-1},\overline{y}_{S,n-1})+k^{Ext}\Sigma_{S},$ (258)
where $\underline{y}_{W,n-1}$ and $\overline{y}_{W,n-1}$ are as above and $\underline{y}_{S,n-1}=\underline{s}_{n-1}-\mu_{S}(t_{n-1})$ and $\overline{y}_{S,n-1}=\overline{s}_{n-1}-\mu_{S}(t_{n-1})$. Note that we set $\pazocal{X}_{0}=\pazocal{X}^{Ref}_{0}$. Given the family of time-varying state spaces $\pazocal{X}_{n}$ we chose equidistant grid points in each of the intervals $[r_{\text{min}},r_{\text{max}}],~{}[w_{\min,n},w_{\max,n}]$ and $[s_{\min,n},s_{\max,n}]$. This leads to the discretized state spaces $\widetilde{\pazocal{X}}_{n}\subset\mathbb{R}^{3}$ used for the calculations in BDP. In our specific case we set $k^{Ref}=3$ and $k^{Ext}=4$.

### v3

@@ H.1 Construction of Running Costs
with $I_{1}=\int_{\mathbb{R}}\tfrac{\Sigma_{S}v+m^{S}}{\sqrt{2\pi}\sqrt{1-\rho^{2}}}\exp\big\{-\tfrac{(v-\rho u)^{2}}{2(1-\rho^{2})}\big\}\mathrm{d}v$. For solving integral $I_{1}$, we note that the function in it is the density of a random variable $V\sim\pazocal{N}(\rho u,1-\rho^{2})$, therefore
$I_{1}$ $=\int_{\mathbb{R}}\tfrac{\Sigma_{S}v}{\sqrt{2\pi}\sqrt{1-\rho^{2}}}\exp\big\{-\tfrac{(v-\rho u)^{2}}{2(1-\rho^{2})}\big\}\mathrm{d}v+\int_{\mathbb{R}}\tfrac{m^{S}}{\sqrt{2\pi}\sqrt{1-\rho^{2}}}\exp\big\{-\tfrac{(v-\rho u)^{2}}{2(1-\rho^{2})}\big\}\mathrm{d}v$
@@
Going back to the calculation of $E(t,k,R)$ and substituting the above, we get
$E(t,k,R)$ $=\int_{u(a)}^{u(b)}\tfrac{1}{\sqrt{2\pi}}e^{k(\Sigma_{W}u+m^{W})}e^{-\tfrac{u^{2}}{2}}(\Sigma_{S}\rho u+m^{S})\mathrm{d}u$
$=\int_{u(a)}^{u(b)}\tfrac{\Sigma_{S}\rho u+m^{S}}{\sqrt{2\pi}}\exp\big\{-\tfrac{u^{2}-2k\Sigma_{W}u+(k\Sigma_{W})^{2}-(k\Sigma_{W})^{2}}{2}+km^{W}\big\}\mathrm{d}u$
$=\int_{u(a)}^{u(b)}\tfrac{\Sigma_{S}\rho u+m^{S}}{\sqrt{2\pi}}\exp\big\{-\tfrac{(u-k\Sigma_{W})^{2}}{2}+\tfrac{(k\Sigma_{W})^{2}}{2}+km^{W}\big\}\mathrm{d}u$
@@
with $I_{2}=\int_{u(a)}^{u(b)}\tfrac{\Sigma_{S}\rho u+m^{S}}{\sqrt{2\pi}}\exp\big\{-\tfrac{(u-k\Sigma_{W})^{2}}{2}\big\}\mathrm{d}u$. Analogous to $I_{1}$, the function within $I_{2}$ is the density of the random variable $U\sim\pazocal{N}(k\Sigma_{W},1)$, which yields
$I_{2}$ $=\Sigma_{S}\rho\int_{u(a)}^{u(b)}\tfrac{u}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u+m^{S}\int_{u(a)}^{u(b)}\tfrac{1}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u$
@@
$I_{3}$ $=\Sigma_{S}\rho\int_{u(a)}^{u(b)}\tfrac{u-k\Sigma_{W}+k\Sigma_{W}}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u$
$=\Sigma_{S}\rho\bigg(\int_{u(a)}^{u(b)}\tfrac{u-k\Sigma_{W}}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u+k\Sigma_{W}\int_{u(a)}^{u(b)}\tfrac{1}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u\bigg)$
$=\Sigma_{S}\rho\bigg(\tfrac{1}{\sqrt{2\pi}}(e^{-\tfrac{(u(a)-k\Sigma_{W})^{2}}{2}}-e^{-\tfrac{(u(b)-k\Sigma_{W})^{2}}{2}})$
$\hskip 8.5359pt+k\Sigma_{W}[\Phi(u(b)-k\Sigma_{W})-\Phi(u(a)-k\Sigma_{W})]\bigg),$ (181)
where $\int_{u(a)}^{u(b)}\tfrac{1}{\sqrt{2\pi}}e^{-\tfrac{(u-k\Sigma_{W})^{2}}{2}}\mathrm{d}u=\Phi(u(b)-k\Sigma_{W})-\Phi(u(a)-k\Sigma_{W})$. The last equality follows from the substitution $x=-\tfrac{(u-k\Sigma_{W})^{2}}{2}$, as well as by solving the integral. Substituting the expression for $I_{3}$ into $I_{2}$ and back into $E(t,k,R)$ gives
$\begin{split}E(t,k,R)&=e^{\tfrac{(k\Sigma_{W})^{2}}{2}+km^{W}(t,y_{W})}\bigg[\tfrac{\Sigma_{S}\rho}{\sqrt{2\pi}}(\exp\big\{-\tfrac{(u(a)-k\Sigma_{W})^{2}}{2}\big\}-\exp\{-\tfrac{(u(b)-k\Sigma_{W})^{2}}{2}\big\})\\ &\hskip 8.5359pt+(\Sigma_{S}\rho k\Sigma_{W}+m^{S}(t,y_{W},y_{S}))[\Phi(u(b)-k\Sigma_{W})-\Phi(u(a)-k\Sigma_{W})]\bigg],\end{split}$ (182)
with $u(w)=\frac{\log(w)-m^{W}(t,y_{W})}{\Sigma_{W}}$. The closed-form expression for $E_{0}(t,k,R)$ is derived from the expression of $E(t,k,R)$ by letting $m^{S}\equiv 1$ and $\Sigma_{S}\equiv 0$, i.e, price is constant with $S(t)=1$ for all $t\in[t_{n},t_{n+1})$.
@@
The computation of the cost functionals requires the evaluation of time dependent integrals, see (170) and (171). However, there are no closed-form expressions to these integrals and therefore approximation is needed. Built-in integral solvers could be used for this task. These solvers usually achieve high accuracy, but have a slow calculation time. Since both, BDP and Q-learning, make repeated use of the cost functionals, a fast alternative to the integral solver needs to be found, which is comparably accurate but faster in calculation. In our study, we use Gaussian quadrature Quarteroni2010 with Legendre polynomials to approximate the integral expressions for $C^{1/2}$ in (170) and (171). For this, let $t_{0},..,t_{k^{I}}$ be $k^{I}+1$ given distinct interpolation points in the interval $t_{n},t_{n+1}]$. The resulting approximation is given by
@@
with $\Delta t=t_{n+1}-t_{n}$ and weights $\omega_{l},~l=1,\ldots,k^{I}$. We find that $k^{I}=2$ is sufficient to obtain good results. Figure 13 shows the solution to $C_{n}^{2}$ obtained by an integral solver and the relative error induced by this Gauss-Legendre quadrature rule. It can be seen that the relative error is less than $2\%$, while acceleration the calculation by a factor of $100$ for the evaluation of the running cost.
Figure 13: Left: The cost functional $C_{n}^{2}$ calculated by an integral solver for $P^{\text{H}}_{n}=3500$, with respect to wind speed and electricity price is shown. Right: The relative error between the Gauss-Legendre quadrature for $k^{I}=2$ and integral solver is shown.
H.2 Construction of Terminal Costs with Penalization and Liquidation
In this work, we consider terminal costs as defined in (31). This means, if the terminal TES temperature $R_{N}=r$ is below or above a critical temperature $r_{\text{crit}}\in[r_{\text{min}},r_{\text{max}}]$, a penalty or reward with respect to the difference $r_{\text{crit}}-r$ is applied. The penalty can be interpreted such that the TES has to be charged to the critical value, thus additional costs are incurring. Conversely, the reward results from the liquidation of the energy saved. This is the energy that would be needed to charge the TES from $r_{\text{crit}}$ to the larger temperature $r$. In both cases, we assign a value to the energy associated with this difference given by a function $g_{\text{Ter}}$. In turn these are multiplied by predetermined fixed prices $s_{\text{Pen}},s_{\text{Liq}}\geq 0$, respectively, to evaluate the TES filling level. Usually, the fixed prices satisfy $s_{\text{Pen}}\geq s_{\text{Liq}}$. In the following, we will describe the penalization and liquidation process in order to construct $g_{\text{Ter}}$.
For the case $r_{\text{crit}}>r$, a penalization cost is applied, which is described by the following scenario. Since the TES temperature is below the critical threshold it needs to be charged manually. We assume that for all $t>t_{\text{E}}$ the produced wind power $P^{\text{W}}(t)=0$ and that the HTHP has to operate at maximum compressor shaft speed $d_{\text{max}}$. Therefore, the TES is charged as fast as possible and the energy consumption is covered solely by grid energy. From (13), the maximum energy consumption $P_{\text{max}}$ in this case is given by
@@ H.2 Construction of Terminal Costs with Penalization and Liquidation
Moreover, the maximum compressor shaft speed $d_{\text{max}}$, yields the maximum outlet temperature
@@
The HTHP run on this continuous-time setting until time $t_{\text{Ter}}=t_{\text{E}}+h(r,r_{\text{crit}})$, for which the TES temperature is raised to $r_{\text{crit}}$. Here, $h(r,r_{\text{crit}})>0$ describes the charging period length, which depends on the current TES state $r$ and the critical value $r_{\text{crit}}$. From the storage dynamics (3), we have
$r_{\text{crit}}=R(t_{\text{Pen}})=r+h(r,r_{\text{crit}})\frac{n_{\text{H}}\dot{m}c_{\text{p,f}}}{m_{\text{s}}c_{\text{p,s}}}(\tau^{\text{HT,out}}_{\text{max}}-T^{\text{SG,in}}),$ (186)
and determine the length of the charging period according to
$h(r,r_{\text{crit}})=\frac{m_{\text{s}}c_{\text{p,s}}}{n_{\text{H}}\dot{m}c_{\text{p,f}}(\tau^{\text{HT,out}}_{\text{max}}-T^{\text{SG,in}})}(r_{\text{crit}}-r).$ (187)
The consumed electrical energy associated is given by the function
$g_{\text{Ter}}(r,r_{\text{crit}})=h(r,r_{\text{crit}})P_{\text{max}}.$ (188)
The liquidation in the case $r_{\text{crit}}<r$ results from the saved energy that would be required to charge the TES from the lower critical temperature $r_{\text{crit}}$ to the higher TES temperature $r$. Analogously to the penalization, we need to charge for a period of length $h(r_{\text{crit}},r)$, with only difference that the role of $r$ and $r_{\text{crit}}$ have changed. Thus, the energy savings associated with the liquidation is given by
$g_{\text{Ter}}(r_{\text{crit}},r)=h(r_{\text{crit}},r)P_{\text{max}}.$ (189)
@@ Appendix I Details State Space Discretization
In order to calculate and store the value function with BDP, the state space $\pazocal{X}$ is going to be discretized into grid points. Moreover, we want to use the additive structure and closed-form solutions of the Ornstein-Uhlenbeck processes (16), to construct a family of discretizations $\pazocal{X}_{n},~n=0,\ldots,N$. By this construction we make sure, that the value function is calculated in regions of interested, i.e., states that are located around the seasonalities. In the first step of the construction, we define the reference sets $\pazocal{X}^{\text{Ref}}_{n}\subset\pazocal{X}$
$\pazocal{X}^{\text{Ref}}_{n}=[r_{\text{min}},r_{\text{max}}]\times[\underline{w}_{n},\overline{w}_{n}]\times[\underline{s}_{n},\overline{s}_{n}],$ (191)
with boundary values $\underline{w}_{n},\overline{w}_{n}\in(0,\infty)$ and $\underline{s}_{n},\overline{s}_{n}\in(-\infty,\infty)$. For our specific case we construct these boundaries with the $k^{\text{Ref}}$-$\sigma$-rule motivated by the normal distribution of the closed-form solutions (26), leading to
$\begin{split}\underline{w}_{n}&=\exp(\mu_{W}(t_{n})-k^{\text{Ref}}\Sigma_{W}),~\overline{w}_{n}=\exp(\mu_{W}(t_{n})+k^{Ref}\Sigma_{W}),\\ \underline{s}_{n}&=\mu_{S}(t_{n})-k^{\text{Ref}}\Sigma_{S},~\overline{s}_{n}=\mu_{S}(t_{n})+k^{\text{Ref}}\Sigma_{S}.\end{split}$ (192)
For states in $\pazocal{X}^{\text{Ref}}_{n}$, we want to achieve a good approximation, meaning that the errors introduced by the simplifications, namely discretization of states and actions as well as quantization of the expected value, should be small. The error of the state and action space discretization can be reduced by choosing a finer grid, i.e., increasing the number of grid points. However, handling the error introduced by the approximation of the expected value is more difficult. A main cause comes from the quantizer $\pazocal{Z}$. For $x\in\pazocal{X}^{\text{Ref}}_{n}$, the quantization points $\pazocal{T}_{n}(x,a,z)$ with $z\in\pazocal{Z}$ are unlikely to be contained in the discrete set $\widetilde{\pazocal{X}}_{n+1}$. It is therefore necessary to interpolate or extrapolate the value function for these states, whereby not only interpolation errors but also extrapolation errors must be taken into account. In order to reduce extrapolation of the value function for states in the reference set, we enlarge them by choosing boundary values $w_{\min,n}\leq\underline{w}_{n},~\overline{w}_{n}\leq w_{\max,n}$ and $s_{\min,n}\leq\underline{s}_{n},~\overline{s}_{n}\leq s_{\max,n}$. This means, we construct an enlarged set $\pazocal{X}_{n}$ with $\pazocal{X}^{\text{Ref}}_{n}\subset\pazocal{X}_{n}$ given by
@@
in the sense that for (almost) every $x\in\pazocal{X}^{\text{Ref}}_{n-1}$, the transition $\pazocal{T}_{n-1}(x,a,z)$ will be contained in $\pazocal{X}_{n}$. This helps to avoid or reduce the need of extrapolation for states in the reference set. The construction of the values for $w_{\min,n},w_{\max,n},s_{\min,n}$ and $s_{\max,n}$ is again motivated by the normal distribution of the closed-form solution (26). For $w_{\min,n}$ and $w_{\max,n}$, we apply the $k^{\text{Ext}}$-$\sigma$-rule leading to
$w_{\min,n}=\exp(m^{W}_{n}(\underline{y}_{W,n-1})-k^{\text{Ext}}\Sigma_{W}),\quad w_{\max,n}=\exp(m^{W}_{n}(\overline{y}_{W,n-1})+k^{\text{Ext}}\Sigma_{W}),$ (194)
for $\underline{y}_{W,n-1}=\log(\underline{w}_{n-1})-\mu_{W}(t_{n-1})$ and $\overline{y}_{W,n-1}=\log(\overline{w}_{n-1})-\mu_{W}(t_{n-1})$. In order to construct $s_{\min,n}$ and $s_{\max,n}$ we proceed analogously. Considering that the electricity price and the wind speed are negatively correlated, i.e., prices are high if wind speeds are low and visa versa, we end up with the following values
$s_{\min,n}=m^{S}_{n}(\overline{y}_{W,n-1},\underline{y}_{S,n-1})-k^{\text{Ext}}\Sigma_{S},\quad s_{\max,n}=m^{S}_{n}(\underline{y}_{W,n-1},\overline{y}_{S,n-1})+k^{\text{Ext}}\Sigma_{S},$ (195)
where $\underline{y}_{W,n-1}$ and $\overline{y}_{W,n-1}$ are as above and $\underline{y}_{S,n-1}=\underline{s}_{n-1}-\mu_{S}(t_{n-1})$ and $\overline{y}_{S,n-1}=\overline{s}_{n-1}-\mu_{S}(t_{n-1})$. Note that we set $\pazocal{X}_{0}=\pazocal{X}^{\text{Ref}}_{0}$. Given the family of time-varying state spaces $\pazocal{X}_{n}$ we chose equidistant grid points in each of the intervals $r_{\text{min}},r_{\text{max}}],~[w_{\min,n},w_{\max,n}]$ and $s_{\min,n},s_{\max,n}]$. This leads to the discretized state spaces $\widetilde{\pazocal{X}}_{n}\subset\mathbb{R}^{3}$ used for the calculations in BDP. In our specific case we set $k^{\text{Ref}}=3$ and $k^{\text{Ext}}=4$.